OpenStudy (anonymous):

At a party, everyone shook hands with everybody else. There were 66 handshakes. How many people were at the party?

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Trick question. 66 * 2 will be the handshakes..

OpenStudy (anonymous):

NVm

OpenStudy (anonymous):

no 66 were the handshakes

OpenStudy (anonymous):

Hold on a minute

OpenStudy (anonymous):

Are you asking these for fun? Or are they part of your school?

OpenStudy (anonymous):

do 66/2

OpenStudy (anonymous):

actually i would say both

OpenStudy (anonymous):

Suppose there were 'n' people. Then each shook hands with (n-1) people. So total number of handshakes = n(n-1)/2 [Why divided by 2? The answer is whether person A shakes hand with person B or person B shakes hand with person A is immaterial as these two situations refer to a single handshake.] Given total no. of handshakes = 66 So n(n-1)/2 = 66 or n^2-n = 66*2 = 132 or n^2-n-132 = 0 or n^2-12n+11n-132 = 0 or n(n-12)+11(n-12) = 0 or (n-12)(n+11) = 0 Therefore either n = 12 or n = -11. But n being no. of people cannot be negative so n = 12

OpenStudy (anonymous):

you see, my math teacher gives us a point for every homework done all together as a class.

OpenStudy (anonymous):

when we get 10 we get a party

OpenStudy (anonymous):

o yea i remember those..

OpenStudy (anonymous):

he said if we get 50 in a row, he'll take us to a field trip all to ourselves wherever we want it

OpenStudy (anonymous):

we have 49 so today he gave him this online work

OpenStudy (anonymous):

thank you 12 In general, with n+1 people, the number of handshakes is the sum of the first n consecutive numbers: 1+2+3+ ... + n. Since this sum is n(n+1)/2, we need to solve the equation n(n+1)/2 = 66. This is the quadratic equation n2+ n -132 = 0. Solving for n, we obtain 11 as the answer and deduce that there were 12 people at the party. Since 66 is a relatively small number, you can also solve this problem with a hand calculator. Add 1 + 2 = + 3 = +... etc. until the total is 66. The last number that you entered (11) is n.

OpenStudy (anonymous):

sorry you were wrong